The width of a rectangle is four feet less than the length. The perimeter is 16 ft. Find the width.
Thanks

Answer by rfer(10432)

w=x=2
l=x+4=6
------------
p=2w+2l
2x+2(x+4)=16
2x+2x+8=16
4x=8
x=2